# An Extensive Study on D2C: Overfitting Remediation in Deep Learning Using a Decentralized Approach

### v1

@@ VI-B Analyzing Loss & Accuracy Curves
Figure 14: Validation accuracy curves for the traditional & the D2C method for different numbers of subsets on AG News.
@@
The loss and accuracy curves generated for AG News depict by far the clearest picture in terms of the improvements brought by the proposed method over the traditional monolithic approach. The loss curve from Figure 13 shows the trend we observed in almost all other cases: dividing the training data into more and more subsets delays and minimizes the increase in validation loss as the training progresses. The convergence though, slows down a bit in this case too. More importantly, with no subsets, the validation loss shoots up sharply after around 8 epochs, which indicates overfitting and impacts the accuracy too as the training progresses as we can see from Figure 14. However, when we use D2C, the sharp increase in the validation loss is mitigated drastically. Even with only two subsets, the improvement is quite significant. When the number of subsets is increased further to more than two, the loss curve becomes quite stable, almost completely eliminating the increasing trend and showing consistent improvement as the training progresses. The accuracy curve (Figure 14) also shows clear indications that D2C improves over the traditional approach. With the traditional approach, the accuracy too starts deteriorating after around the 14th epoch, indicating severe overfitting, an issue that no longer exists when we use D2C for any value of $N$ or $E$.
@@
The experiments on all these datasets from 3 different datasets prove that using the D2C method can indeed address overfitting and control the abrupt increase of true error in the form of maintaining a lower validation loss during training. This also shows that D2C works well in different domains and a diverse set of tasks.
VI-C Accuracy, F1 Score, AUC-ROC, and Log Loss
@@
We evaluated our proposed D2C on test sets using Accuracy, F1 Score, Log Loss, and AUC-ROC to capture a comprehensive view of model performance in terms of correctness, class balance, prediction confidence, and discrimination ability. Reporting these final metrics on test sets was essential to confirm that improvements generalized well beyond the training and validation data, verifying true effectiveness in unseen scenarios. We used the average from multiple repetitive runs (to test for repeatability) for each of these metrics. We used the provided test sets with each of the datasets for proper comparisons. In the tables, the top row represents the traditional centralized approach and the rest of the rows represent different variations of the D2C approach. The bold entities represent the best performance in the tables.
@@
Image Classification: Let’s take a look at the performance metrics from the benchmark image classification datasets we primarily used using different settings (the traditional approach and the D2C method with different values of $N$ and $E$).
@@
TABLE II: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on CIFAR-10.
@@
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.8596 0.8591 0.7346 0.98697
N=2, E=1 0.8611 0.8611 0.7042 0.98693
N=3, E=1 0.8613 0.8609 0.6682 0.98704
N=4, E=1 0.8586 0.8583 0.6516 0.98690
N=5, E=1 0.8552 0.8551 0.6389 0.98628
N=7, E=1 0.8539 0.8535 0.6198 0.98640
N=9, E=1 0.8531 0.8525 0.6068 0.98605
@@
TABLE III: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on Fashion MNIST.
@@
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.9314 0.9315 0.3467 0.99559
N=2, E=1 0.9323 0.9319 0.3460 0.99549
N=3, E=1 0.9330 0.9329 0.3179 0.99545
N=3, E=2 0.9319 0.9316 0.3303 0.99552
N=3, E=3 0.9337 0.9337 0.3225 0.99590
N=5, E=1 0.9322 0.9321 0.2692 0.99589
N=7, E=1 0.9338 0.9334 0.2467 0.99584
N=7, E=2 0.9305 0.9304 0.2575 0.99600
N=9, E=1 0.9323 0.9323 0.2341 0.99600
@@
TABLE IV: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on MNIST.
@@
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.9952 0.9952 0.0258 0.99995
N=2, E=1 0.9957 0.9956 0.0271 0.99995
N=3, E=1 0.9958 0.9958 0.0240 0.99998
N=3, E=2 0.9951 0.9950 0.0250 0.99996
N=5, E=1 0.9957 0.9956 0.0202 0.99998
N=7, E=1 0.9951 0.9950 0.0215 0.99997
N=9, E=1 0.9957 0.9957 0.0179 0.99998
N=11, E=1 0.9959 0.9959 0.0161 0.99999
N=11, E=2 0.9958 0.9958 0.0170 0.99998
@@
As we can see from Table II, Table III, and Table IV, the best accuracy, MCC and AUC-ROC were achieved using the D2C method for all three datasets. For CIFAR-10, the number of subsets, $N$, was $3$ for the best model in terms of accuracy. The number of epochs before each round of central averaging, $E$, was just $1$. This approach showed an improvement with an accuracy of $86.13\%$ over the traditional approach with no subsets, which yielded an accuracy of $85.96\%$. Initially, the accuracy improved as we increased the number of subsets, but suffered a consistent drop after it was further increased past $3$. As we discussed, dividing the training set into too many subsets may impact training negatively due to a lack of representative samples. The experimental outcomes show further evidence of that in this case.
@@
For Fashion MNIST too, the best accuracy, MCC, and AUC-ROC were achieved using our proposed Divide2Conquer method. The number of subsets, $N$, was $7$ in this case for the best model in terms of accuracy. The number of epochs before each round of central averaging, $E$, was once again just $1$. This approach again showed an improvement with an accuracy of $93.38\%$ over the traditional approach with no subsets, which yielded an accuracy of $93.14\%$. An increasing trend is visible in accuracy as we increase the number of subsets till it reaches the best conditions with $N=7$.
@@
In the case of MNIST too, the best accuracy, MCC, and AUC-ROC were achieved using our proposed method. The number of subsets, $N$, was $11$ in this case for the best model. The number of epochs before each round of central averaging, $E$, was $1$. The improvement of $0.07$ in error rate over the traditional approach is significant for the MNIST dataset. In this case, the least amount of log loss is also achieved with the best model in terms of accuracy.
@@
We can also see that as we increase the number of subsets, $N$, there is a clear trend of decreasing log loss. As we discussed earlier, increasing the number of subsets further minimizes the effect of noise and outliers. The decrease in log loss is a result of that phenomenon. Even though the best model isn’t usually selected based on log loss values, in the case of similar accuracy (for example, the accuracies for $N=3,E=3$ and $N=7,E=1$ in Fashion MNIST dataset), the best model can be selected based on log loss. Log Loss evaluates how well the model predicts probabilities, penalizing confident but incorrect predictions, which is critical for understanding model reliability. A lower log loss means a more robust and confident model, so the model with a lower log loss is preferred when the test accuracy or F1 score is similar.
@@
The comparison of subset numbers in the best-performing models across CIFAR-10, Fashion MNIST, and MNIST provides critical insights. For CIFAR-10, the optimal model utilized only $3$ training subsets, while the best models for Fashion MNIST and MNIST required $7$ and $11$ subsets, respectively. Notably, increasing the subset count beyond $3$ led to declining accuracy for CIFAR-10—a trend not observed with the other datasets. This discrepancy is likely due to the distinct characteristics of each dataset. While all contain $10$ classes, CIFAR-10, a smaller dataset with RGB images and complex class representations, presents a more challenging classification task compared to the grayscale images of Fashion MNIST and MNIST. Consequently, effective training on CIFAR-10 demands larger sample sizes per subset, and when these subsets become too small, performance degrades. In contrast, Fashion MNIST and MNIST, identical in dataset size, differ in task complexity, with Fashion MNIST presenting a slightly more challenging classification task than MNIST. This complexity results in the optimal Fashion MNIST model requiring fewer subsets than MNIST, as each subset’s sample size needs to be sufficient for adequate training.
@@
This analysis suggests a key principle for applying the D2C method effectively: to divide training data into more subsets, datasets must be large enough to allow each subset ample training data. Larger datasets benefit more from divided training, requiring more subsets to achieve optimal performance.
@@
TABLE V: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on FER2013.
@@
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subsets 0.6498 0.6420 2.0112 0.89934
N=2, E=1 0.6511 0.6394 1.9101 0.89875
N=2, E=2 0.6567 0.6444 1.8450 0.9032
N=2, E=3 0.6490 0.6389 1.8718 0.8975
N=3, E=1 0.6491 0.6356 1.6540 0.90104
N=3, E=2 0.6538 0.6465 1.6990 0.89870
N=3, E=3 0.6539 0.6428 1.7315 0.9002
N=5, E=1 0.6500 0.6368 1.5157 0.90153
N=7, E=1 0.6521 0.6432 1.4821 0.90148
@@
FER2013, a challenging dataset for facial expression recognition, represents a task that is particularly prone to overfitting. As seen in Table V, our proposed D2C method outperformed the traditional approach and achieved the highest accuracy, MCC, and AUC-ROC, with the best-performing model using only 2 subsets ($N=2$) and 2 training epochs before each central averaging step ($E=2$). D2C yielded a notable accuracy improvement, reaching 65.67% compared to 64.98% in the traditional model without subsets. Unlike previous datasets, where a single epoch ($E=1$) was optimal, FER2013 benefited from 2 epochs, likely due to the dataset’s complexity, requiring additional training time to capture meaningful features before model averaging.
@@
A decreasing trend in log loss was observed with more subsets, though no consistent pattern emerged with varying epochs ($E$). Additionally, due to FER2013’s class imbalance, accuracy alone may not reliably indicate the best model, so, the F1-score, a metric considering class-wise precision, provides a valuable perspective. Interestingly, the optimal F1-score model (3 subsets, $E=2$) differs from the accuracy-optimized model, with D2C achieving an F1 score of 0.6465 versus 0.642 in the baseline model. This F1-score improvement underscores D2C’s effectiveness in addressing overfitting, particularly in class-imbalanced, complex tasks.
Audio Classification: Now let us take a look at the performance metrics from the Combined Audio dataset.
@@
TABLE VI: Performance Metrics with Traditional Approach vs. Different Settings Using D2C on the Combined Audio Dataset.
@@
Model Settings Accuracy F1-Score Log Loss AUC-ROC
No Subset 0.6953 0.6870 1.3419 0.96570
N=2, E=1 0.6942 0.6879 1.0811 0.96553
N=2, E=2 0.6908 0.6819 1.1636 0.9632
N=3, E=1 0.6920 0.6909 1.2032 0.96416
N=3, E=2 0.6874 0.6830 1.0033 0.96438
N=4, E=1 0.6921 0.6804 1.2840 0.96214
@@
We can see from Table VI that in the Combined Audio dataset, the highest accuracy was achieved using the traditional training approach rather than our proposed D2C method. The traditional model attained a 69.53% accuracy, slightly outperforming the D2C approach, which reached 69.42% with $2$ subsets. Accuracy declined further as the number of subsets increased, likely due to the smaller dataset size (11,632 samples), limiting effective data availability per subset. As a result, dividing the dataset into smaller portions likely impacted performance by reducing the sample count for each class, a significant challenge given the dataset’s complexity and diversity, as it combines samples from multiple sources.
@@
However, despite marginally lower accuracy, the D2C method showed reduced log loss, indicating a trade-off between accuracy and confidence in predictions. In cases of comparable accuracy, lower log loss can be preferable as it indicates more calibrated predictions. Moreover, given the dataset’s class imbalance, the F1-score is a more reliable metric here. The D2C approach yielded a slight improvement in the F1-score (0.6909 vs. 0.687), indicating better generalization across classes. This suggests that D2C supports robust classification with fewer, more distinctive errors, especially in class-imbalanced, complex tasks like this one. In summary, while D2C did not maximize accuracy, it achieved comparable F1 performance with lower log loss, resulting in a model that generalizes well and offers more confident predictions in challenging, class-imbalanced datasets. However, dividing the dataset into too many subsets does worsen the performance of the models in this case.
@@
This analysis leads us to the conclusion that to apply the Divide2Conquer method effectively, the dataset needs to be sufficiently large so that each of the subsets can have enough training data. It is however noteworthy that even with a comparatively smaller dataset, our proposed method still shows some performance improvement in some metrics.
@@
TABLE VII: Performance Metrics with Traditional Approach vs. Different Settings Using the D2C Method on AG News.

### v2

@@ 5.2. Analyzing Loss & Accuracy Curves
Figure 9. Validation accuracy curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on Fashion MNIST.
However, the loss curve isn’t the only important factor during the training process. The relation between log loss and accuracy may not be so linear. That is why it is important to observe the validation accuracy too. As we can see from Figure 9, for all the different values of the number of subsets, $N$, even though the loss reaches its minimum much earlier, the accuracy shows an increasing trend (even if slightly) almost throughout the entire training process, right up until the 60th epoch. This shows why continuing training for a significant period might be necessary even if the validation loss starts increasing. This also shows why early stopping based on loss curves may not be a good idea in both cases: the traditional approach and the DUA-D2C method. This is something we observed in the cases of all these datasets. Also, we can notice a visible improvement in terms of validation accuracy in Figure 9. At the very least, dividing the training set into multiple subsets ensures that the maximum accuracy is attained while keeping the log loss relatively in control, which characterizes a more confident and robust model with better generalization ability.
Let us also take a look at the validation loss curves generated using FER2013 while keeping the number of subsets, $N$, constant and varying the number of local epochs before each round of central aggregation, $E$. We varied $E$ between $1$ to $3$ and Kept $N$ constant at $2$ and $3$ separately.
Figure 10. Validation loss curves for different numbers of epochs before each round of central averaging, while keeping the number of subsets, N = 2 & 3.
We can see from Figure 10 that when we vary the number of epochs before each round of central aggregation, $E$, there is no clear indication of any kind of change in the trend or the rate of change in the validation losses as the training progresses. This is unlike when we varied the number of subsets, $N$, keeping $E$ constant, where we saw that increasing the number of subsets delayed and minimized the increase in validation loss as the training progressed quite clearly. Also, the convergence got slower. But here, for both $N=2$ (shades of blue) and $N=3$ (shades of green), the trends from the validation curves are quite similar for different values of $E$ throughout the training phase. In light of the above observations, we can conclude that varying the number of epochs before each round of central aggregation, $E$ does not have any obvious implications on log loss as we saw in the case of varying the number of subsets, $N$. However, an increased number of local epochs before each round of central aggregation means more scope for learning from each of the subsets before combining and assigning the weights, sometimes even overfitting on a particular subset a bit more. We will observe the effect of varying the number of epochs further in the later subsections when we observe the final accuracy and log loss in each case. The implication of varying $E$ is likely to be different on different datasets from different domains as we will see in the case of the number of subsets too.
Audio Classification: Overfitting is a common problem in audio classification tasks. But it is more apparent when the model is trained on a particular dataset, and then tested on data from a different source. One way of tackling this issue is by combining datasets to increase the diversity in both training and test/validation data. We combined the datasets TESS (dupuis2010toronto, 23), CREMA-D (cao2014crema, 4), and RAVDESS (livingstone2018ryerson, 18) for this task. Let us take a look at the validation loss and accuracy curves generated using this Combined Audio dataset.
Figure 11. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets, N, on the Combined Audio dataset.
As we can see from Figure 11, there is overfitting to some extent in all cases, which causes the validation loss to creep up after a few epochs but the rate at which this phenomenon takes place is much less when the training set is divided into multiple subsets. Once again, the rate of increase in validation loss goes down as we increase the number of subsets. Even though we converted the 1D time series audio files to 2D Log-Mel spectrograms to form something like an image, the classification based on these Mel-specs differs greatly from traditional image classification. So, the overfitting-resistant nature of models using our proposed method being apparent in this case is a very promising sign. However, the small size of this dataset may not be ideal in the context of applying DUA-D2C. We will analyze it in detail in the next subsection based on several performance metrics.
Text Classification: Text classification is an important case study when it comes to testing the DUA-D2C method. For both image classification and audio classification, we used 2D data and CNN to carry out the tasks. But text classification deals with 1D sequences, and as we mentioned in the previous section, we used bi-directional LSTM in this case. We used AG News (zhang2015character, 33), a benchmark dataset to test our method primarily in the NLP domain. Let us observe the validation loss and accuracy curves generated using the AG News dataset first.
Figure 12. Validation loss curves for the traditional & the DUA-D2C method for different numbers of subsets on AG News.
Figure 13. Validation accuracy curves for the traditional & the DUA-D2C method for different numbers of subsets on AG News.
@@
The loss and accuracy curves generated for AG News depict by far the clearest picture in terms of the improvements brought by the proposed method over the traditional monolithic approach. The loss curve from Figure 12 shows the trend we observed in almost all other cases: dividing the training data into more and more subsets delays and minimizes the increase in validation loss as the training progresses. The convergence though, slows down a bit in this case too. More importantly, with no subsets, the validation loss shoots up sharply after around 8 epochs, which indicates overfitting and impacts the accuracy too as the training progresses as we can see from Figure 13. However, when we use DUA-D2C, the sharp increase in the validation loss is mitigated drastically. Even with only two subsets, the improvement is quite significant. When the number of subsets is increased further to more than two, the loss curve becomes quite stable, almost completely eliminating the increasing trend and showing consistent improvement as the training progresses. The accuracy curve (Figure 13) also shows clear indications that DUA-D2C improves over the traditional approach. With the traditional approach, the accuracy too starts deteriorating after around the 14th epoch, indicating severe overfitting, an issue that no longer exists when we use DUA-D2C for any value of $N$ or $E$.
@@
The experiments on all these datasets from 3 different datasets prove that using the DUA-D2C method can indeed address overfitting and control the abrupt increase of true error in the form of maintaining a lower validation loss during training. This also shows that DUA-D2C works well in different domains and a diverse set of tasks.
5.3. Accuracy, F1 Score, AUC-ROC, Cohen’s Cappa, Log Loss and MCC
@@
We evaluated our proposed DUA-D2C on test sets using Accuracy, F1 Score, AUC-ROC, Cohen’s Cappa, Log Loss, and Matthew’s Correlation Coefficient (MCC) to capture a comprehensive view of model performance in terms of correctness, class balance, prediction confidence, and discrimination ability. Reporting these final metrics on test sets was essential to confirm that improvements generalized well beyond the training and validation data, verifying true effectiveness in unseen scenarios. We used the average from multiple repetitive runs (to test for repeatability) for each of these metrics. We used the provided test sets with each of the datasets for proper comparisons. In the tables, the top row represents the traditional centralized approach, and the rest of the rows represent different variations of the DUA-D2C approach. The bold entities represent the best performance in the tables.
@@
Image Classification: Let’s take a look at the performance metrics from the benchmark image classification datasets we primarily used using different settings (the traditional approach and the DUA-D2C method with different values of $N$ and $E$).
@@
Table 2. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on CIFAR-10.
@@
Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
No Subsets 0.8604 0.8601 0.98731 0.8458 0.7198 0.8458
N=2, E=1 0.8626 0.8628 0.98704 0.8476 0.7002 0.8476
N=3, E=1 0.8652 0.8648 0.98761 0.8501 0.6491 0.8501
N=3, E=2 0.8630 0.8624 0.98719 0.8478 0.6215 0.8478
N=4, E=1 0.8595 0.8591 0.98710 0.8439 0.6416 0.8439
N=5, E=1 0.8614 0.8613 0.98676 0.8461 0.6347 0.8461
N=7, E=1 0.8556 0.8553 0.98658 0.8393 0.6126 0.8398
N=9, E=1 0.8548 0.8542 0.98664 0.8385 0.5876 0.8390
@@
Table 3. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on Fashion MNIST.
@@
Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
No Subsets 0.9312 0.9313 0.99546 0.9237 0.3492 0.9238
N=3, E=1 0.9338 0.9336 0.99570 0.9264 0.2865 0.9265
N=3, E=2 0.9344 0.9340 0.99571 0.9271 0.3006 0.9272
N=3, E=3 0.9332 0.9329 0.99580 0.9258 0.3008 0.9258
N=5, E=1 0.9327 0.9323 0.99581 0.9252 0.2618 0.9253
N=7, E=1 0.9321 0.9321 0.99591 0.9246 0.2383 0.9246
N=7, E=2 0.9322 0.9319 0.99600 0.9247 0.2377 0.9247
N=9, E=1 0.9300 0.9299 0.99590 0.9222 0.2314 0.9222
@@
Table 4. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on MNIST.
@@
Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
No Subsets 0.9952 0.9952 0.99995 0.9947 0.0258 0.9947
N=2, E=1 0.9954 0.9954 0.99991 0.9949 0.0294 0.9949
N=3, E=1 0.9955 0.9955 0.99995 0.9950 0.0225 0.9950
N=3, E=2 0.9953 0.9953 0.99995 0.9948 0.0259 0.9948
N=5, E=1 0.9957 0.9957 0.99998 0.9952 0.0188 0.9952
N=7, E=1 0.9960 0.9959 0.99998 0.9956 0.0183 0.9956
N=9, E=1 0.9963 0.9963 0.99999 0.9959 0.0148 0.9959
N=11, E=1 0.9953 0.9953 0.99998 0.9948 0.0158 0.9948
N=11, E=2 0.9959 0.9959 0.99998 0.9954 0.0156 0.9954
@@
As we can see from Table 2, Table 3, and Table 4, the best accuracy, F1-score, MCC, Cohen’s Cappa, log loss, and AUC-ROC were achieved using the DUA-D2C method for all three datasets. For CIFAR-10, the number of subsets, $N$, was $3$ for the best model in terms of accuracy. The number of epochs before each round of central aggregation, $E$, was just $1$. This approach showed an improvement with an accuracy of $86.52\%$ over the traditional approach with no subsets, which yielded an accuracy of $86.04\%$. Initially, the accuracy improved as we increased the number of subsets, but suffered a decreasing trend after it was further increased past $3$. As we discussed, dividing the training set into too many subsets may impact training negatively due to a lack of representative samples. The experimental outcomes show further evidence of that in this case.
@@
For Fashion MNIST too, the best metrics were achieved using our proposed DUA-D2C method. The number of subsets, $N$, was once again $3$ in this case for the best model in terms of accuracy. The number of epochs before each round of central aggregation, $E$, was $2$ in this case. This approach again showed an improvement with an accuracy of $93.44\%$ over the traditional approach with no subsets, which yielded an accuracy of $93.12\%$. An increasing trend is visible in ROC-AUC as we increase the number of subsets till it reaches the best conditions with $N=7$.
@@
In the case of MNIST too, the best metrics were all achieved using our proposed method. The number of subsets, $N$, was $9$ in this case for the best model. The number of epochs before each round of central aggregation, $E$, was $1$. The improvement of $0.11$ in error rate over the traditional approach is very significant for the MNIST dataset. In this case, the least amount of log loss is also achieved with the best model in terms of accuracy.
@@
We can also see that as we increase the number of subsets, $N$, there is a clear trend of decreasing log loss. As we discussed earlier, increasing the number of subsets further minimizes the effect of noise and outliers. The decrease in log loss is a result of that phenomenon. Even though the best model isn’t usually selected based on log loss values, in the case of similar accuracy, the best model can be selected based on log loss. Log Loss evaluates how well the model predicts probabilities, penalizing confident but incorrect predictions, which is critical for understanding model reliability. A lower log loss means a more robust and confident model, so the model with a lower log loss is preferred when the test accuracy or F1 score is similar.
@@
The comparison of subset numbers in the best-performing models across CIFAR-10, Fashion MNIST, and MNIST provides critical insights. For CIFAR-10, the optimal model utilized only $3$ training subsets, while the best models for Fashion MNIST and MNIST required 33 (77 in terms of ROC-AUC) and $9$ subsets, respectively. Notably, increasing the subset count beyond $3$ led to significantly declining accuracy for CIFAR-10, a trend not observed with the other datasets. This discrepancy is likely due to the distinct characteristics of each dataset. While all contain $10$ classes, CIFAR-10, a smaller dataset with RGB images and complex class representations, presents a more challenging classification task compared to the grayscale images of Fashion MNIST and MNIST. Consequently, effective training on CIFAR-10 demands larger sample sizes per subset, and when these subsets become too small, performance degrades. In contrast, Fashion MNIST and MNIST, identical in dataset size, differ in task complexity, with Fashion MNIST presenting a slightly more challenging classification task than MNIST. This complexity results in the optimal Fashion MNIST model requiring fewer subsets than MNIST, as each subset’s sample size needs to be sufficient for adequate training.
@@
This analysis reiterates a key principle for applying the DUA-D2C method effectively: to divide training data into more subsets, datasets must be large enough to allow each subset enough training data. Larger datasets benefit more from divided training, requiring more subsets to achieve optimal performance.
@@
Table 5. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on FER2013.
@@
Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
No Subsets 0.6498 0.6420 0.89934 0.5756 2.0112 0.5759
N=2, E=1 0.6512 0.6457 0.90262 0.5781 1.6723 0.5783
N=2, E=2 0.6510 0.6381 0.90360 0.5776 1.5595 0.5777
N=2, E=3 0.6480 0.6338 0.89610 0.5740 1.6505 0.5750
N=3, E=1 0.6471 0.6365 0.89910 0.5735 1.6408 0.5739
N=3, E=2 0.6482 0.6346 0.90088 0.5743 1.6311 0.5749
N=3, E=3 0.6420 0.6232 0.89670 0.5671 1.4744 0.5682
N=5, E=1 0.6537 0.6370 0.90528 0.5816 1.1934 0.5818
N=5, E=2 0.6482 0.6308 0.90020 0.5741 1.4750 0.5747
N=7, E=1 0.6474 0.6339 0.90244 0.5732 1.3299 0.5743
@@
FER2013, a challenging dataset for facial expression recognition, represents a task that is particularly prone to overfitting. As seen in Table 5, our proposed method outperformed the traditional approach and achieved the best accuracy, Cohen’s Cappa, MCC, log loss, and AUC-ROC, with the best-performing model using 5 subsets ($N=5$) and 1 training epoch before each central aggregation step ($E=1$). DUA-D2C yielded a notable accuracy improvement, reaching 65.37% compared to 64.98% of the traditional model without subsets. However, D2C performed even better with up to 65.67% of accuracy using the best settings. A decreasing trend in log loss was observed with more subsets, though no consistent pattern emerged with varying epochs ($E$).
@@
FER2013 is a particularly important test case for us as this dataset represents a comparatively imbalanced class distribution. Due to FER2013’s class imbalance, accuracy alone may not reliably indicate the best model, so, the F1-score, a metric considering class-wise precision, provides a valuable perspective. Interestingly, the optimal F1-score model (2 subsets, $E=1$) differs from the accuracy-optimized model, with D2C achieving an F1 score of 0.6457 versus 0.642 in the baseline model. This F1-score improvement underscores DUA-D2C’s effectiveness in addressing overfitting, particularly in class-imbalanced, complex tasks.
Audio Classification: Now let us take a look at the performance metrics from the Combined Audio dataset. This is another important case as combining datasets like TESS, CREMA-D, and RAVDESS is likely to lead to Non-IID data, and this one is another imbalanced and smaller dataset.
@@
Table 6. Performance Metrics with Traditional Approach vs. Different Settings Using DUA-D2C on the Combined Audio Dataset.
@@
Model Settings Accuracy F1-Score ROC-AUC Cohen Kappa Log Loss Matthews CC
No Subsets 0.6953 0.6870 0.96570 0.6684 1.3419 0.6692
N=2, E=1 0.6869 0.6787 0.96226 0.6592 1.7225 0.6601
N=2, E=2 0.7010 0.6973 0.96360 0.6747 1.5021 0.6751
N=3, E=1 0.6946 0.6849 0.96284 0.6678 1.4576 0.6681
N=3, E=2 0.6775 0.6695 0.96560 0.6492 1.0571 0.6504
N=4, E=1 0.6930 0.6805 0.96423 0.6659 1.3829 0.6662
N=5, E=1 0.6938 0.6862 0.96523 0.6668 1.2160 0.6673
@@
We can see from Table 6 that in the Combined Audio dataset, the best metrics were achieved using our proposed DUA-D2C method once again. The traditional model attained a 69.53% accuracy, while the DUA-D2C approach, reached 70.10% with $2$ subsets and $E=2$. Accuracy declined as the number of subsets increased, likely due to the smaller dataset size (11,632 samples), limiting effective data availability per subset. As a result, dividing the dataset into smaller portions likely impacted performance by reducing the sample count for each class, a significant challenge given the dataset’s complexity and diversity, as it combines samples from multiple sources.
@@
However, despite declining accuracy, the DUA-D2C method showed reduced log loss as $N$ increased to 3, indicating a trade-off between accuracy and confidence in predictions. Given the dataset’s class imbalance, the F1-score is a more reliable metric here. The DUA-D2C approach yielded a significant improvement in the F1-score (0.6973 vs. 0.687), indicating better generalization across classes. This suggests that DUA-D2C supports robust classification with fewer, more distinctive errors, especially in class-imbalanced, complex tasks like this one. However, the traditional approach outperformed both DUA-D2C and D2C in terms of ROC-AUC for this dataset. In summary, while DUA-D2C did not maximize ROC-AUC, it achieved better metrics overall, resulting in a model that generalizes well and offers more confident predictions in challenging, class-imbalanced datasets. However, dividing the dataset into too many subsets does worsen the performance of the models in this case.
@@
This analysis leads us to the conclusion that to apply the DUA-D2C method effectively, the subsets need to be sufficiently large so that each of them can have enough training data. It is however noteworthy that even with a comparatively smaller dataset, DUA-D2C still shows performance improvement in most metrics.
@@
Table 7. Performance Metrics with Traditional Approach vs. Different Settings Using the DUA-D2C Method on AG News.